# Puffy Venuses: the Mass - Radius Impact of Carbon-Rich Atmospheres on Lava Worlds

### v1

@@ 1 Introduction
Observing outgassed secondary atmospheres offers an unparalleled window to understanding rocky exoplanet interiors and habitability. With the arrival of JWST, detecting such atmospheres on rocky exoplanets becomes feasible (Ostberg & Kane, 2019). Recent JWST phase-curve observation of the ultrashort period, underdense super-Earth 55 Cancri e indicate the exoplanet hosts a CO/CO2 atmosphere (Hu et al., 2024). The emergence of a population of such underdense, highly irradiated rocky planets (Piette et al., 2023) motivates the question: can thick CO/CO2 dominated atmospheres explain these planets’ low density? What is the maximum extent of these Venus-like atmospheres, and can they create mass-radius relationships distinct from that of Earth-like planets?
@@
Despite some caveats and unknowns, it is not difficult to envision a carbon-rich formation pathway for rocky exoplanets via pebble accretion, which efficiently grows planets from primitive material (Ormel & Klahr, 2010; Johansen & Lambrechts, 2017; Johansen et al., 2021; Izidoro et al., 2021). A planet may receive volatile-rich pebbles drifting from beyond the snow line (Sato et al., 2016; Ida et al., 2019). Alternatively, it may form beyond the snow line rapidly, before the dispersal of the protoplanetary disc, and subsequently migrate inwards to become hotter - and thus puffier, making them easier to detect via transits (Lambrechts et al., 2019; Izidoro et al., 2021). Potential caveats - mechanisms of carbon loss - include the atmospheric recycling with the surrounding disc (Johansen et al., 2021), the dissolution and burial of carbon into the core (Fischer et al., 2020; Johansen et al., 2023), as well as the destruction of carbonaceous grains in the disc through photolysis (Lee et al., 2010; Gail & Trieloff, 2017; Binkert & Birnstiel, 2023). However, previous works on these mechanisms are primarily in the context of volatile delivery during Earth formation, which is under lively debate (see, e.g., Johansen et al. (2021)). Therefore, their impact on the exoplanet population remains unclear.
The above discussions suggest that carbon-rich exoplanets may exist. Yet current exoplanet interior models have overlooked the potential significance of a thick carbon atmosphere in favor of hydrogen or water-rich layers (e.g., Lacedelli et al. (2022)). In this study, we seek to understand the maximum impact of a carbon-dominated atmosphere on the radius of rocky exoplanets, a scenario we term puffy Venuses. We do this by constructing coupled atmosphere-magma ocean (MO) models since the thickest atmospheres would result in strong insulation and, thus, likely necessitate hot interiors where the mantle is substantially molten. We incorporate C-H-O thermochemistry, building on recent advances in modeling lava worlds in and beyond the Solar System (Sossi et al., 2020; Bower et al., 2022; Gaillard et al., 2022). We demonstrate strong radius inflation among irradiated, low-mass puffy Venuses. For instance, a 0.16 - 0.3 $R_{\oplus}$ carbon-dominated atmosphere can exist on an Earth-mass, highly-irradiated ($T_{eq}\in[1500,2000]$K) planet with modest (ordinary chondrite, 1200 ppm (Schaefer & Fegley, 2017)) carbon abundances. We identify nine puffy Venus candidates and highlight our cases for TOI-561 b and 55 Cnc e.
Figure 1: Three sample atmospheric profiles of Earth-mass planets with OC total C and H abundances. a (fiducial): Atmospheric composition profile of a planet with solar C/O = 0.55 and $T_{eq}$ = 1500 K. The shaded region denotes the convective layer. b: Same as a but with a more reducing atmosphere of C/O = 0.8. c: Same as a but with $T_{eq}=2000$ K. d: The atmosphere pressure-temperature of planets a (solid green), b (dash-dot green) and c (solid orange). For comparison, the dashed orange line is an atmosphere at $T_{eq}=2000$ K that stays convective to the MO surface.
Both a higher $T_{eq}$ and a more reduced environment promote higher surface temperatures, resulting in larger atmospheres. For reference, $z_{atm}$= 0.162, 0.233, 0.307 $R_{\oplus}$ for planets a, b and c, respectively.
Figure 2: Mass - radius relations of puffy Venuses. a: M-R relations compared to observed exoplanets. The grey line is the M-R for terrestrial planets with an Earth-like core mass fraction of 0.325 (Plotnykov & Valencia, 2020). Colored solid lines are puffy Venus radii assuming different $T_{eq}$, with the fiducial, OC-like volatile C and H abundances. The orange dashed line is the M-R for $T_{eq}=2000$ K planets with a convective deep atmosphere to the MO surface, instead of a deep radiative zone. The shaded region is the range of M-R space an air-less rocky world can occupy for all possible Fe/Si $\in[0,1]$. Planets less dense than this region require a volatile component. Measured exoplanet samples are from the NASA Exoplanet Archive (2023), and only contain those with relative errors below 30%. The nine planets highlighted with black circles are our preferred puffy Venus candidates. In rising mass order, they are GJ 1252 b, TOI-500 b, TOI-561 b, Kepler-10 b, Kepler-36 b, HD 219134 b, HD 3167 b, WASP-47 e, and 55 Cnc e. TOI-561 b’s radius is from Patel et al. (2023); its mass is from Brinkman et al. (2023) b: M-R of puffy Venuses with a range of chondritic C abundances, more opaque lines correspond to less C. They are color-coded by temperature - $T_{eq}$ = 1000 K (blue) and 2000 K (orange). The dashed orange line is for wet CI composition, while the solid line with the same color is for dry CI composition. The dotted portion of dry CI, 1000 K line are atmospheres that should experience graphite precipitation; see text. c: M-R of models assuming different chemical scenarios - I (solid), local chemical equilibrium throughout the atmosphere, II (dash-dot), chemical equilibrium at the MO surface, and a CO2 + H2O atmosphere (dotted). Fiducial OC C and H abundances are used. Same color-coding as b. The solid 1000 K and 2000 K lines in a, the 2 OC lines in b, and the 2 Scenario I lines in c are the same.
@@ 2 Magma Ocean - Atmosphere model
We model planets with a carbon-rich atmosphere in thermochemical equilibrium with a magma ocean and obtain the height of the atmosphere $z_{atm}$. We distribute a total volatile C and H budget between a homogeneous MO and a 1-D, semi-grey radiative-convective atmosphere, which allows for the development of deep radiative zones that Selsis et al. (2023) demonstrated to be important for thick secondary atmospheres. We find $z_{atm}$ as the height where the atmosphere thermal radiation reaches a transit optical depth of 2/3.
@@
We consider an atmosphere composed of 6 species - CO, CO2, CH4, H2, H2O and O2. We recognize that besides the volatile C and H species, these exoplanets can also contain refractory C and H species not available to the atmosphere, but their abundance is likely limited (see Appendix B.2). The atmosphere is either (I) in chemical equilibrium (CE) at each layer and in CE with the MO at the interface, or (II) well-mixed with a composition set by the CE with the magma ocean at the interface. Case I represents the scenario when chemical kinetics is faster than vertical mixing, while case II represents the reverse. We also consider (III) CO2-H2O atmospheres as an end member case. In all scenarios, the MO and the atmosphere are coupled at their interface via solubility laws. We find the mass of the MO by calculating the adiabatic mantle thermal profile assuming an Earth-like lower-mantle petrology.
Our model is extensively described in Appendix A.

### v2

@@ 1 Introduction
Observing outgassed secondary atmospheres offers an unparalleled window to understanding rocky exoplanet interiors and habitability. With the arrival of JWST, detecting such atmospheres on rocky exoplanets becomes feasible (Ostberg & Kane, 2019). Recent JWST phase-curve observation of the ultrashort period, underdense super-Earth 55 Cancri e indicate the exoplanet hosts a CO/CO2 atmosphere (Hu et al., 2024). The emergence of a population of such underdense, highly irradiated rocky planets (Piette et al., 2023) motivates the question: can thick CO/CO2 dominated atmospheres explain these planets’ low density? What is the maximum extent of these Venus-like atmospheres, and can they create mass - radius (M-R) relationships distinct from that of Earth-like planets?
@@
Despite some caveats and unknowns, it is not difficult to envision a carbon-rich formation pathway for rocky exoplanets via pebble accretion, which efficiently grows planets from primitive material (Ormel & Klahr, 2010; Johansen & Lambrechts, 2017; Johansen et al., 2021; Izidoro et al., 2021), thus avoiding some volatile loss pathways associated with hierarchical accretion (Gu et al., 2024). A planet may receive volatile-rich pebbles drifting from beyond the snow line (Sato et al., 2016; Ida et al., 2019). Alternatively, it may form beyond the snow line rapidly, before the dispersal of the protoplanetary disc, and subsequently migrate inwards to become hotter - and thus puffier, making them easier to detect via transits (Lambrechts et al., 2019; Izidoro et al., 2021). Potential caveats - mechanisms of carbon loss - include the atmospheric recycling with the surrounding disc (Johansen et al., 2021), the dissolution and burial of carbon into the core (Fischer et al., 2020; Johansen et al., 2023), as well as the destruction of carbonaceous grains in the disc through photolysis (Lee et al., 2010; Gail & Trieloff, 2017; Binkert & Birnstiel, 2023). However, previous works on these mechanisms are primarily in the context of volatile delivery during Earth formation, which is under lively debate (see, e.g., Johansen et al. (2021)). Therefore, their impact on the exoplanet population remains unclear.
[…]
@@ 2 Magma Ocean - Atmosphere model
We model planets with a carbon-rich atmosphere in thermochemical equilibrium with a magma ocean and obtain the height of the atmosphere $z_{atm}$. We distribute a C and H total volatile budget between a homogeneous MO and the atmosphere modeled via a 1-D semi-grey radiative-convective treatment (Section 2.4), which allows for the development of deep radiative zones that Selsis et al. (2023) demonstrated to be important for thick secondary atmospheres. We find $z_{atm}$ as the height where the atmosphere thermal radiation reaches a transit optical depth of 0.56 (de Wit & Seager, 2013; Heng & Kitzmann, 2017).
@@
We consider an atmosphere composed of 6 species - CO, CO2, CH4, H2, H2O, and O2 (Section 2.1). We recognize that besides the volatile C and H species, these exoplanets can also contain C and H species not available to the atmosphere, such as those dissolved into the metallic core. But their abundance is likely limited (Section 4.3.2). We consider the atmosphere to be either (I) in chemical equilibrium (CE) at each layer and in CE with the MO at the interface or (II) well-mixed with a composition set by the CE with the magma ocean at the interface. Case I represents the scenario when chemical kinetics is faster than vertical mixing, while case II represents the reverse. We also consider (III) pure CO2-H2O atmospheres as an end member case. In all scenarios, the MO and the atmosphere are coupled at their interface via solubility laws (see Section 2.4). We find the mass of the MO by calculating the adiabatic mantle thermal profile assuming an Earth-like lower-mantle petrology.
2.1 Atmospheric chemical equilibrium & equations of state
We use three gas-phase chemical reactions to couple the six volatile species in our atmosphere model - $\mathrm{H_{2}}$, $\mathrm{O_{2}}$, $\mathrm{H_{2}O}$, $\mathrm{CO_{2}}$, $\mathrm{CO}$ and $\mathrm{CH_{4}}$:
$\mathrm{CO+0.5\,O_{2}}\Leftrightarrow\mathrm{CO_{2}}$ (1)
$\mathrm{H_{2}+0.5\,O_{2}}\Leftrightarrow\mathrm{H_{2}O}$ (2)
$\mathrm{CO_{2}+2\,H_{2}}\Leftrightarrow\mathrm{CH_{4}+O_{2}}$ (3)
As detailed below, the equilibrium molar fractions of these gases, $y_{i}$’s, are solved given total pressure (P), temperature (T), and atomic ratios C/O and C/H. In the fiducial Scenario I, we use the local P and T throughout the radiative-convective atmosphere, which is obtained from the numerical integration of the structure equations (see Section 2.2). We ensure convergence given that the molar fraction of gases feed back into the structure equations via density and mean specific heat capacity. In Scenario II, instead, we only solve for CE $y_{i}$’s at the magma ocean surface P-T. The chemical equilibrium between the atmosphere and the magma ocean is handled by the Henrian solubility laws using $y_{i}$’s and the surface pressure, see Section 2.4.
Each reaction corresponds to an equilibrium constant $K_{eq}$:
[…]
where $f_{i}$ denotes the fugacity of gas species $i$. Fugacity describes the availability of a compound for chemical reactions. We calculate $K_{eq}(T)$ utilizing the thermochemical data from the NIST database (Thermodynamics Research Center, 2021). As numerical fits to the Shomate equation, the NIST data are accurate up to 6000 K. We linearly extrapolate the NIST data beyond 6000 K, but we note that our fiducial models predict deep atmosphere temperatures $\lesssim 7000$ K (see fig. 1d). We note that higher temperature behaviors of $K_{eq}$’s remain qualitatively sensible.
In general, a real gas $i$’s fugacity is a function of both the local pressure and temperature, and the composition of the gas mixture. In other words, $f_{i}=f(T,P,y_{i},y_{j\neq i})$, where $y_{i}=N_{i}/N$ is the molar fraction. In practice, comparable atmosphere speciation models often assume ideal gas, $f_{i}=y_{i}P$ (Bower et al., 2022; Shorttle et al., 2024), or non-ideal gases in ideal mixture $f_{i}=f(T,P,y_{i})$ (e.g., Sun & Lee 2022) to simplify the calculation, with the notable exception of Tian & Heng (2024), who incorporated non-ideal mixing for a subset of their species. We choose to adopt the non-ideal gas in ideal mixture approach to balance realism and tractability. We discuss this choice further in Section 4.3.4.
Under our assumption, fugacity and local P-T are related by
$f_{i}=\phi_{i,pure}y_{i}P,$ (7)
where $\phi_{i,pure}$ is the fugacity coefficient calculated from the equation of state (EoS) of pure gas $i$. Here, P is the total pressure at a radial location in the atmosphere. We use the modified Lee and Kesler EoS developed by Duan et al. (1992) for CO2 and H2O, and use a generalized version of it for the other 4 species (Duan et al., 1996). The Duan et al. (1992) EoS is accurate to 1273 K and 8000 bars and approximately accurate for higher P-T, while the Duan et al. (1996) EoS is accurate to 2000K and 25000 bars. The equation of state is expressed in terms of compressibility $Z_{i}$
$Z_{i}(P,T)\equiv\frac{PV}{R_{gas}T},$ (8)
where $V$ is the molar volume and $R_{gas}$ is the gas constant. Then from first principles, $\phi_{i,pure}$ can be found via (Duan et al., 1992)
$\ln\phi_{i,pure}(P,T)=\int_{0}^{P}(Z(P,T)-1)\frac{\mathrm{d}P^{\prime}}{P^{\prime}}.$ (9)
In practice, we use the EoS and the analytical $\phi(Z,P,T)$ relationship from Duan et al. (1992) to calculate $Z_{i}$ and $\phi_{i}$ tables. We then interpolate from these tables during atmosphere integration. We assume constant atomic C/O and C/H at each layer of the atmosphere
$\mathrm{C/O}=\frac{y_{\mathrm{CO_{2}}}+y_{\mathrm{CO}}+y_{\mathrm{CH_{4}}}}{2y_{\mathrm{CO_{2}}}+y_{\mathrm{CO}}+2y_{\mathrm{O_{2}}}+y_{\mathrm{H_{2}O}}},$ (10)
$\mathrm{C/H}=\frac{y_{\mathrm{CO_{2}}}+y_{\mathrm{CO}}+y_{\mathrm{CH_{4}}}}{4y_{\mathrm{CH4}}+2y_{\mathrm{H_{2}O}}+2y_{\mathrm{H_{2}}}}$ (11)
Finally, we close the system by enforcing $\sum y_{i}=1$, and solve 4 through 11 numerically for the local gas mixture in equilibrium: $y_{i}(P,T,\mathrm{C/O},\mathrm{C/H})$.
The C/O and C/H ratios are set by the MO-atmosphere interaction. C/O is a proxy for the atmosphere’s redox state. Because the mantle has vastly more redox power than the outgassed atmosphere, the former should control the latter’s redox state. In the literature, the mantle redox state is usually parameterized by mineral buffers, most commonly the iron-wustite buffer (e.g., Gaillard et al. 2022; Bower et al. 2022; Tian & Heng 2024). However, these buffers are experimentally calibrated to Earth mantle P-T, below $\sim$ 3000K (Ballhaus et al., 1991; Hirschmann, 2021). This is easily exceeded in our deep atmospheres, which reach $>$6000 K for the most irradiated planets. We, therefore, choose to specify the C/O for our atmospheres as an input variable.
Although this approach effectively divorces the redox of the atmosphere from mantle geochemistry, it has merits beyond necessity. Firstly, how redox reactions such as IW, active in the deep mantle, influence the oxidation state at the surface of the magma ocean is complex (Hirschmann, 2012, 2022; Gaillard et al., 2022). Further, the extent to which extrasolar magma oceans share similar composition and geochemistry as their terrestrial counterparts is uncertain. In fact, greater diversity in exoplanet composition has been inferred from polluted white dwarfs (Putirka & Xu, 2021). Moreover, the C/O ratio of an exoplanet atmosphere is potentially observable in the JWST era and has started to be constrained for some hot Jupiters (Line et al., 2021; Changeat et al., 2022) and sub-Neptunes (Benneke et al., 2024).
To summarize, chemical speciation is calculated via solving 4 through 11, using P, T, C/O and C/H as input. For Scenario I, integrating atmospheric profile provides local $P(r)$ and $T(r)$, while Scenario II only requires MO surface P and T. C/O is a model input, while C/H is iteratively found via the MO-atmosphere mass balance that we describe in Section 2.4.
2.2 1-D radiative-convective atmosphere profile
We developed a simple 1-D radiative-convective model based on the semi-gray treatment of Guillot (2010), hereafter G10. This scheme treats the planet’s thermal emission and stellar irradiation separately, each characterized by a mean opacity. This scheme allows us to account for a deep radiative zone that a long-living lava-world atmosphere likely develops (Selsis et al., 2023).
We choose this treatment over a full-physics radiative-convective atmosphere model both for the simple model’s conceptual clarity and for necessity. Since our model is concerned with constraining the atmospheric thickness for the population of C-rich exoplanets rather than predicting in detail the spectral signatures of specific exoplanets, our simple treatment suffices. Moreover, while recent efforts (e.g., Gordon et al. 2022) have expanded our understanding of the relevant molecules’ radiative properties up to P $\leq$ 1000 bars and T $\lesssim$ 4000 K, the deep atmospheres we probe (T $\gtrsim$ 3000 K, P $\gtrsim$ 3000 bars) remains elusive. Moreover, the cross-species (e.g., CO-CO2, CO2-O2) collision-induced absorption is not well-characterized to high temperatures (see review in Chubb et al. 2024).
Overall, our model integrates the atmospheric pressure-temperature profile using the differential form of G10’s system of equations for isotropic irradiation and switching to a dry, non-ideal adiabat when appropriate. We also account for atmosphere self-gravity. For Scenario I, local gas composition $y_{i}$ is informed by thermochemical equilibrium. $y_{i}$ then feeds back to the adiabatic lapse rate and density profile. For Scenario II, the atmospheric profile takes constant $y_{i}$ throughout the atmosphere. See Section 2.4 for details.
2.2.1 Radiative Temperature Gradient
We first derive the radiative temperature gradient following G10. Here we summarize our treatment. We start with the first 3 Eddington moments of the intensity $I_{\nu\mu}$, defined as -
$(J_{\nu},H_{\nu},K_{\nu})=\frac{1}{2}\int_{-1}^{1}I_{\nu\mu}(1,\mu,\mu^{2})\mathrm{d}\mu$ (12)
where $J_{\nu}$, $4\pi H_{\nu}$, and $4\pi K_{\nu}/c$ corresponds to the energy of the beam, the radiation flux, and radiation pressure, respectively. $\mu=\cos\theta$ is the incident angle. Assuming the incoming visible, and outgoing thermal radiation fields are independent, their wavelength-integrated moments become
$(J_{v},H_{v},K_{v})\equiv\int_{visible}(J_{\nu},H_{\nu},K_{\nu})\mathrm{d}\nu$ (13)
$(J_{th},H_{th},K_{th})\equiv\int_{thermal}(J_{\nu},H_{\nu},K_{\nu})\mathrm{d}\nu$ (14)
In a planar atmosphere in thermodynamic equilibrium, these moments of radiation are related by
$\frac{\mathrm{d}H_{v}}{\mathrm{d}m}=\kappa_{v}J_{v}$ (15)
$\frac{\mathrm{d}K_{v}}{\mathrm{d}m}=\kappa_{v}H_{v}$ (16)
$\frac{\mathrm{d}H_{th}}{\mathrm{d}m}=\kappa_{th}(J_{th}-B)$ (17)
$\frac{\mathrm{d}K_{th}}{\mathrm{d}m}=\kappa_{th}H_{th},$ (18)
where $B=\sigma/\pi T^{4}$ is the local black-body radiance, and $\sigma$ is the Stefan-Boltzmann constant. $m=-\int\rho\mathrm{d}r$ is the atmosphere column mass, integrating from the “top” of the atmosphere downwards. The local mean gas density $\rho$ at the radial location $r$ is found using the non-ideal EoS
$\rho=\frac{P}{R_{gas}T}\frac{\sum y_{i}\mu_{i}}{\sum y_{i}Z_{i}},$ (19)
where $\mu_{i}$ is the molecular mass of species $i$. Because $Z\geq 1$ in our P-T range, our non-ideal atmosphere tends to be less dense and puffier than their ideal counterparts.
$\kappa_{v}$ and $\kappa_{th}$ are mean opacities
$\kappa_{v}=J_{v}^{-1}\int_{v}\kappa_{\nu}J_{\nu}\mathrm{d}\nu$ (20)
$\kappa_{th}=J_{th}^{-1}\int_{th}\kappa_{\nu}J_{\nu}\mathrm{d}\nu$ (21)
$\kappa_{v}$ and $\kappa_{th}$ are complex functions of atmospheric composition and stellar irradiation. Our model allows for any functional form of visible and thermal absorption coefficients $\kappa_{v,th}(P,T)$. But we, by default, use constant $\kappa$’s. We choose conservative values that fit the P-T observations of the Venusian atmosphere: $\kappa_{v}=7.524\cdot 10^{-6}\mathrm{cm^{2}/g}$ and $\kappa_{th}=3.3\cdot 10^{-3}\mathrm{cm^{2}/g}$. We also tested $\kappa\propto P$, which is motivated by the pressure-broadening of collision-induced absorption lines (Heng et al., 2012; Tolento & Robinson, 2019). The two cases produce similar results, so we choose the simpler constant case. We discuss the caveats of this treatment and show that our main result is robust against our opacity choices in Section 4.3.1.
We can then define optical depths for both radiation fields
$\mathrm{d}\tau_{v,th}=\kappa_{v,th}\mathrm{d}m$ (22)
Local thermodynamic equilibrium dictates
$0=\kappa_{th}(J_{th}-B)+\kappa_{v}J_{v}$ (23)
This is equivalent to writing $H_{v}+H_{th}\equiv H_{int}$, where $H_{int}=\sigma T_{int}^{4}/(4\pi)$ is the outgoing heat flux at the bottom of the atmosphere, or the internal heat flux [^1] We assume no surface effect as our atmosphere is thick. Geophysically speaking, $H_{int}$ roots from the residual heat of planet formation and the radiogenic heating of the planet, encoding its thermal history. We choose $T_{int}=0$ K as the fiducial value to model a planet at thermal equilibrium, yielding the least amount of inflation with no additional heat warming up the deep atmosphere. Choosing Earth or Jupiter-level internal heating only marginally inflates the atmosphere since the high stellar irradiation dominates a puffy Venus’s energy budget.
To close the system of equations, we invoke the Eddington approximation following G10
$K_{th}=\frac{1}{3}J_{th}$ (24)
$K_{v}=\frac{1}{3}J_{v}$ (25)
They are strictly correct for an isotropic atmosphere or in the two-stream approximation and are generally accurate to the first order of magnitude (Parmentier & Guillot, 2014). For the incoming radiation field, combining 15,16, and 25 gives
$(J_{v}(\tau_{v}),H_{v}(\tau_{v}))=(J_{v0},H_{v0})\exp(-\sqrt{3}\tau_{v}),$ (26)
where $J_{v0}=-\sqrt{3}H_{v0}$ are outer boundary conditions at $\tau=0$. $H_{v0}=-\sigma T_{eq}^{4}/(\sqrt{3}\pi)$ is the incoming stellar radiation, which is a downward flux and, thus, negative.
We define the opacity ratio folloiwng G10 as $\gamma\equiv\kappa_{v}/\kappa_{th}$. Rewriting Equation 23 in terms of $J_{th}$ and taking its derivative, we have
$\frac{\mathrm{d}J_{th}}{\mathrm{d}m}=\frac{\mathrm{d}B}{\mathrm{d}m}-\frac{\mathrm{d}}{\mathrm{d}m}(\gamma J_{v})$ (27)
By combining the above with Equations 18 and 24, we obtain
$\kappa_{th}H_{th}=\frac{1}{3}\Big{[}\frac{\mathrm{d}B}{\mathrm{d}m}-\Big{(}\frac{\mathrm{d}\gamma}{\mathrm{d}m}J_{v}+\frac{\mathrm{d}J_{v}}{\mathrm{d}m}\gamma\Big{)}\Big{]}.$ (28)
We express the P and T dependencies for $\gamma$ and T dependency of $B$ explicitly, and use Equations 16 and 25 to rewrite the last term, obtaining
$\begin{split}0=&\frac{1}{3}\frac{\mathrm{d}B}{\mathrm{d}T}\frac{\mathrm{d}T}{\mathrm{d}m}\\ &-\frac{1}{3}\Big{(}\frac{\partial\gamma}{\partial T}\frac{\mathrm{d}T}{\mathrm{d}m}J_{v}+\frac{\partial\gamma}{\partial P}\frac{\mathrm{d}P}{\mathrm{d}m}J_{v}\Big{)}\\ &-(\kappa_{v}\gamma H_{v}+\kappa_{th}H_{th}).\end{split}$ (29)
We enforce hydrostatic equilibrium
$\mathrm{d}P=g(r)\mathrm{d}m=-\rho(r)g(r)\mathrm{d}r,$ (30)
and rearrange Equation 29, using the definition of $B$ and switching variables $\mathrm{d}r=-1/\rho\cdot\mathrm{d}m$ to obtain the radiative temperature gradient
$\begin{split}\frac{\mathrm{d}T}{\mathrm{d}r}\Big{|}_{rad}=-&\rho\Big{[}\frac{\partial\gamma}{\partial P}g(r)J_{v}\\ &+3(\kappa_{v}\gamma H_{v}+\kappa_{th}H_{th})\Big{]}\\ &\times\Big{(}\frac{4\sigma}{\pi}T^{3}-\frac{\partial\gamma}{\partial T}J_{v}\Big{)}^{-1}\end{split}$ (31)
Here $g(r)=GM(r)/r^{2}$ is the local gravitational acceleration accounting for atmosphere self-gravity, and $M(r)$ is the mass of the planet enclosed at radius $r$, which we numerically integrate via
$\mathrm{d}M=4\pi r^{2}\rho(r)\mathrm{d}r,$ (32)
2.2.2 Non-ideal Adiabat
When the radiative gradient exceeds the adiabat, the atmosphere profile follows a non-ideal adiabat - adiabat generalized to real-gas EoS. The adiabat is given by
$\frac{\mathrm{d}T}{\mathrm{d}P}\Big{|}_{ad}=\frac{T}{\overline{c_{p}}}\frac{\mathrm{d}V}{\mathrm{d}T}\Big{|}_{P},$ (33)
where $\overline{c_{p}}$ is the mean specific heat. We calculate $\overline{c_{p}}(y_{i},T)$ via
$\overline{c_{p}}(T)=\sum_{i}y_{i}c_{p,i}(T).$ (34)
Specific heat for individual species, $c_{p,i}(T)$, are calculated using a Shomate equation fit to tabulated data provided by NIST (Thermodynamics Research Center, 2021)
$c_{p,i}(T)=A+B\cdot t+C\cdot t^{2}+D\cdot t^{3}+E\cdot^{-2},$
where $A$, $B$, $C$, $D$ and $E$ are compound-specific fitting constants and $t=T/1000$. We linearly extrapolate $c_{p,i}(T)$ beyond 6000 K. We calculate the volume gradient for an individual gas from the EoS
$\frac{\mathrm{d}V_{i}}{\mathrm{d}T}\Big{|}_{P}=Z_{i}\frac{R_{gas}}{P}+\frac{\mathrm{d}Z_{i}}{\mathrm{d}T}\Big{|}_{P}\frac{R_{gas}T}{P}$ (35)
$\frac{\mathrm{d}V}{\mathrm{d}T}\Big{|}_{P}=\sum y_{i}\frac{\mathrm{d}V_{i}}{\mathrm{d}T}\Big{|}_{P}+\sum V_{i}\frac{\mathrm{d}y_{i}}{\mathrm{d}T}\Big{|}_{P}.$ (36)
2.2.3 Boundary conditions & transit height
We integrate the atmospheric profiles numerically from the top $r_{top}$ to the MO surface at $R_{s}$, using the LSODA integrator in the SciPy package (Virtanen et al., 2020). Specifically, we integrate 30 for $P$, equations 31 and 33 for $T$, 32 for $M(r)$, and 22 for $\tau_{v}$ and $\tau_{th}$. At the top of the atmosphere, we choose $P_{top}=0.1$ Pa, $\tau_{v,top}=10^{-9}$ and $\tau_{th,top}=\tau_{th,top}/\gamma$. We choose fiducial $\gamma=2.28\cdot 10^{-3}$ (see Section 4.3.1). The top temperature is given by Equation 28 of G10, taking their incident angle $\mu_{*}=1/\sqrt{3}$ and their irradiation temperature $T_{irr}^{4}=4T_{eq}^{4}$:
$T_{top}=\Big{[}\frac{1}{2}T_{int}^{4}+\frac{2}{\sqrt{3}}T_{eq}^{4}\Big{(}1+\frac{\sqrt{3}}{2}\gamma\Big{)}\Big{]}$ (37)
We numerically integrate the transit, or chord, optical depth (see G10) in the thermal wavelength $\tau_{th,ch}$, and define the height of the atmosphere $z_{atm}$ as the location where $\tau_{th,ch}=0.56$, a number derived for transiting isothermal atmospheres from first principles (de Wit & Seager, 2013; Heng & Kitzmann, 2017). Since the high atmosphere in G10 is close to isothermal, this is an appropriate choice. Choosing the visible transit location $\tau_{v,ch}=0.56$, or the commonly-used $\tau_{th,ch}=2/3$ as the criteria do not strongly impact our results.
From integrating the atmospheric profile, we also obtain the total mass of the atmosphere $M_{atm}$, as well as the total mass of C and H in the atmosphere $M_{C,H,atm}$ for the mass balance calculation (section 2.4).
2.3 Thermal structure of the mantle & the magma ocean depth
We integrate the adiabatic mantle temperature profile $T(r,\rho)$ using the Vinet EoS for the lower mantle, a mixture of bridgmanite and magnesiowustite (Plotnykov & Valencia, 2020). We discuss the inaccuracies of this simplified approach in Section 4.3.3 and deem them acceptable.
To find the extent of mantle melting, we combine two sets of solidus $T_{sol}(P)$ and liquidus $T_{liq}(P)$ curves. We use peridotite curves from Morschhauser et al. (2011) for the low-pressure end and chondritic curves from Monteux et al. (2016) up to 140 GPa. Beyond that, we linearly interpolate the Monteux et al. (2016) curves. We find the local melt fraction via linear interpolation
$\phi_{melt}=\frac{T-T_{sol}}{T_{liq}-T_{sol}}.$ (38)
2.4 MO-atmosphere C & H mass balance
To account for the chemical exchange between the atmosphere and magma ocean we assume vigorous mixing to bring the atmosphere and the magma ocean to equilibrium at their interface, similar to previous magma ocean models (e.g Bower et al. 2022; Lichtenberg et al. 2021; Gaillard et al. 2022). This means that we set the bulk magma ocean C and H abundances by the chemical equilibrium and solubility relations at the MO surface. We describe the partitioning of the dissolved vs. outgassed gas species with the Henrian solubility relations from Lichtenberg et al. (2021). These relations link the partial pressures of individual species to their dissolved abundances in the MO:
$X_{i}=a\cdot(y_{i,s}P_{s})^{b},$ (39)
where $X_{i}$ is the abundance of dissolved gas $i$ in the MO by weight, and $a$ and $b$ are constant fits to experimental data. $P_{s}$ is the total pressure, and $y_{i,s}$ is the molar fraction, both at the magma ocean surface. As we describe later, these are from the numerically integrated atmospheric profile.
We find the total dissolved elemental C and H mass, $M_{C,\,MO}$ and $M_{H,\,MO}$ via bookkeeping
$M_{C,\,MO}=M_{MO}\sum_{i}k_{i,C}\frac{\mu_{C}}{\mu_{i}}X_{i}$ (40)
$M_{H,\,MO}=M_{MO}\sum_{i}k_{i,H}\frac{\mu_{H}}{\mu_{i}}X_{i},$ (41)
where $k_{i,C}$ and $k_{i,H}$ are the number of C and H atoms in compound $i$. $\mu_{C}$, $\mu_{H}$ and $\mu_{i}$ are molar weights of C, H and species $i$. The total C and H abundances ($x_{C,tot}$, $x_{H,tot}$) then found via
$x_{C,tot}=\frac{1}{M_{tot}}(M_{C,MO}+M_{C,atm}).$ (42)
$x_{H,tot}=\frac{1}{M_{tot}}(M_{H,MO}+M_{H,atm}).$ (43)
Overall, the coupled atmosphere magma ocean model takes $x_{C,tot}$, $x_{H,tot}$, $M_{s}$, $R_{s}$, and atmospheric C/O as inputs. Numerically, we start with guesses of the top-of-atmosphere (ToA) conditions $r_{top}$ and $M_{tot}$, and in Scenario I, the atmospheric C/H. We integrate from $r_{top}$ to the MO surface at $R_{s}$ to find the MO surface conditions $P_{s}$, $y_{i,s}$ and $T_{s}$, as well as $M(r=R_{s})$, the mass enclosed by the MO surface from integration (Sections 2.1 and 2.2). The first three of these then inform the MO volatile abundance $M_{C,H,MO}$ via equations 39 to 41. Together with the atmospheric C and H masses $M_{C,H,atm}$, we find bulk abundances $x_{C,tot}$ and $x_{H,tot}$. We then iterate the ToA $r_{top}$ and $M_{tot}$, as well as atmospheric C/H until we find the atmosphere-MO pair that produces the desired $x_{C,tot}$ and $x_{H,tot}$, as well as the self-consistent requirement that the mass enclosed at the MO surface from atmosphere integration equals the mass of the rocky planet below
$M(r=R_{s})=M_{s}.$ (44)
In the chemically quenched Scenario II, in addition to $M_{tot}$, $r_{top}$ and C/H, the atmosphere model also iterates for the MO surface conditions $P_{s}$ and $T_{s}$ to find the correct gas mixture $y_{i}$ that satisfies the aforementioned requirements. To close the system, we also enforce self-consistency on the atmosphere P-T profile -
$(P(r=R_{s}),T(r=R_{s}))=(P_{s},T_{s}).$ (45)